OpenStudy (anonymous):

simplify 1/x-3+4/x divided by 4/x-1/x-3 is the answer: 5x-12/3(x-4) ?

OpenStudy (anonymous):

for sure

OpenStudy (anonymous):

i wouldn't lie, really

OpenStudy (anonymous):

\[\frac{\frac{1}{x-3}+\frac{4}{x}}{\frac{4}{x}-\frac{1}{x-3}}\] is question?

OpenStudy (anonymous):

im not saying ur laying but u did not simplify

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{x+4(x-3)}{4(x-3)-x}\]

OpenStudy (anonymous):

i already did

OpenStudy (anonymous):

\[\frac{5x-12}{3x-12}\], simplified i get your answer

OpenStudy (anonymous):

the answer is 5x-12/3x-12

OpenStudy (anonymous):

fully simplified leave the denominator as 3(x-4)

OpenStudy (anonymous):

so yes i did 5x-12/3(x-4)

OpenStudy (anonymous):

yea thats right

OpenStudy (anonymous):

ok thank you i wanted to make sure

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Satellite73 i know you gave me the correct answer

OpenStudy (anonymous):

not to be picky but there is nothing "simpler" about \[\frac{5x-12}{3x-12}\] vs \[\frac{5x-12}{3(x-4)}\]

OpenStudy (anonymous):

i now and i did it

OpenStudy (anonymous):

i was only asking look at my question

OpenStudy (anonymous):

np

OpenStudy (anonymous):

(5x - 12) / (3x -12)

OpenStudy (anonymous):

thanks